# Neutrino flavor instabilities in neutron star mergers with moment transport: slow, fast, and collisional modes

### v1

@@ IV.3 Scattering processes and isotropy-breaking modes
where all the definitions can be found in Eq. (42), we do not reproduce them here for brevity. We assumed, without loss of generality, $\Delta_{\mathcal{N}}>\overline{\Delta}_{\mathcal{N}}$ (i.e., $A>0$), and we dropped the indices $\mathcal{N}$ and $\mathcal{F}$ as there is no difference here. Since $|\Delta_{\mathcal{N}}-\overline{\Delta}_{\mathcal{N}}|\geq\Delta_{\mathcal{N}}+\overline{\Delta}_{\mathcal{N}}$ (i.e., $|A|\leq G$), we see that the growth rate of the isotropy-breaking modes is smaller. However, the inclusion of isotropic scattering processes, which enter the flux equation but not the number density one [see Eq. (10)], can lead a priori to $\gamma_{\mathcal{F}}+(G\alpha_{\mathcal{F}})/A>-\gamma_{\mathcal{N}}+(G\alpha_{\mathcal{N}})/A$, if the inevitable increase in $\gamma_{\mathcal{F}}>\gamma_{\mathcal{N}}$ is compensated by an increase $\alpha_{\mathcal{F}}>\alpha_{\mathcal{N}}$.
@@ IV.4 Vacuum term and slow modes
Therefore, the imaginary part of Eq. (18) is not modified. However, in the regions with a gapless or no CFI, slow modes could play a role. We show the growth rates obtained from homogeneous and isotropic LSA in Fig. 9, top panel. They result from a combination of collisional and slow modes. In the bottom panel, we display the ratio of this growth rate with the default isotropic CFI-only (top panel of Fig. 2). As in Fig. 7, we show in turquoise the newly unstable regions, and in pink the regions which were unstable without the vacuum term, but are stable with it. As expected from the argument above, the regions with the largest differences in growth rates correspond to the regions of gapless instability or of stability in Fig. 2. The results of Fig. 9 assume a normal ordering of neutrino masses—the results for the inverted ordering are very similar and we don’t show them for brevity. The newly unstable regions are mostly associated with very small growth rates ($10^{4}\,\mathrm{s}^{-1}$), such that we do not expect a large impact of these slow modes. In particular, local studies of the slow instability in those regions would not be justified, since large-scale advection occurs on comparable or smaller timescales. Nevertheless, as we will show in the following, including the anisotropies of the neutrino field can lead to a significant enhancement of these growth rates (see Sec. V.2).
@@ V Results for anisotropic neutrino distributions
In this section and for the remainder of this paper, we no longer assume that neutrino distributions are isotropic, and we include the actual nonzero fluxes $\mathbf{F}_{\alpha\alpha}\neq\mathbf{0}$. This has two consequences. First, angular crossings between the neutrino and antineutrino distributions can occur, which are connected to fast instabilities [113, 114, 33]. Second, the separation between isotropy-preserving and isotropy-breaking modes—which can be done when the stability matrix has a block structure, see Eq. (40)—is not possible anymore. Furthermore, we also now consider different collision rates for number and flux, $\Gamma_{N}$ and $\Gamma_{F}$.
@@ V.1 Results at $k^{\prime}=0$, no vacuum term
Although our goal here is not to discuss in depth the origin of the angular crossings that lead to FFI, we note that our regions of FFI are consistent with other works. First, the instability patterns (small regions above and below the HMNS, extended regions in the tidal arms) are similar to the ones found in a 5 ms postmerger snapshot of another binary NSM simulation in [29] (see also Appendix D). Another comparison can be made with the recent detailed study [51], which used multiangle Boltzmann neutrino transport on top of fixed fluid backgrounds obtained from a $1.35M_{\odot}-1.35M_{\odot}$ NSM simulation leading to a long-lived ($1.3\,\mathrm{s}$) HMNS. While in the simulation we study the HMNS collapses to a black hole at $8.5\,\mathrm{ms}$ post-merger, some of the same physical mechanisms occur. For instance, the thin layers near the surface of the HMNS correspond to near-isotropic configurations with similar number densities of $\nu_{e}$ and $\bar{\nu}_{e}$, such that a small difference in flux between these two species leads to a crossing (see Sec. III.B.1. in [51]). Ref. [51] points out other mechanisms, mostly based on the larger emission of $\bar{\nu}_{e}$ than $\nu_{e}$ at the surface of the HMNS, and the presence of an optically thick accretion disk with $\mu_{\nu}^{\mathrm{eq}}>0$. However, some of the processes leading to crossings are dependent on having sufficient angular resolution, and may be smeared out by using a two-moment scheme assuming maximum entropy angular distributions.
@@
Since the CFI is subdominant in the regions far from the HMNS, we focus upon the central region, see the right panel of Fig. 10. Contrary to what was observed in CCSNe [47], we find that the CFI is still largely present, extending the regions of instability. The growth rates are significantly smaller, but they could still lead to a build-up of flavor coherence throughout the equatorial region. The light-crossing time of a $1\,\mathrm{km}$ region is $\tau\simeq 1/(3\times 10^{5})\,\mathrm{s}$, which shows that instabilities with growth rates below a few times $10^{5}\,\mathrm{s}^{-1}$ can only have a cumulative, nonlocal effect (contrary to the FFI), which remains to be studied. We also caution that our LSA assumes constant and homogeneous flavor on-diagonal densities, such that our results only give us information on the local properties of the instabilities. Although this is beyond the scope of this work, we refer to [45] for a discussion of a possible way to perform a nonlocal LSA which takes into account the large-scale inhomogeneities of the classical neutrino field.
V.2 Fast, collisional and slow modes: towards the full picture
@@
Note we have assumed $f_{\nu_{x}}=f_{\bar{\nu}_{x}}$. Although this formula may not be quantitatively exact, it ensures a FFI growth rate of order $\sqrt{2}G_{F}\mathcal{N}\sim 10^{10}\,\mathrm{s}^{-1}$—much larger than the CFI growth rates—for all regions where a crossing should be present (assuming maximum entropy angular distributions). We thus superimpose this estimate on the results at $\mathbf{k}^{\prime}=\mathbf{0}$, see Fig. 11. In the left panel, we see that now almost all CFI regions at large distances from the HMNS are overwhelmed by a FFI, but the zoomed-in version (right panel) still shows an extension of the instability regions in thin layers near the surface of the HMNS and, more significantly, in the tidal arms. We note that a detailed comparison of the FFI regions between Figs. 10 and 11 shows that the expression (38) can underestimate the FFI growth rate since the $\mathbf{k}^{\prime}=\mathbf{0}$ mode does not necessarily have the largest growth rate.
@@ V.3 Summary: focus on a few example locations
In order to summarize the several growth rate determinations we have obtained in this work, and how different regions of a NSM simulation are subject to different flavor instabilities, we focus here on four locations identified on Fig. 12. The instability growth rate for these four points is shown in Fig. 13, where the marker matches the one used in Fig. 12, and each column corresponds to a different LSA assumption. From left to right, first assuming isotropic neutrino background distributions, “Iso CFI” is the multi-energy CFI, compared with the energy-averaged “Method A” (19) and “Method B” (21) (see Sec. IV.1), and with various multi-energy extensions: with many-body corrections to the absorption opacities (“Iso CFI, m-b,” Sec. IV.2), including scattering opacities (“Iso CFI + scatt,” Sec. IV.3), or with the vacuum term (“Iso CFI + NO,” Sec. IV.4). Then, we show the results obtained when we take into account the nonzero neutrino fluxes: “Aniso ($\mathbf{k}^{\prime}=\mathbf{0}$)” when the vacuum term is not included (Sec. V.1), “Estimate $\sigma_{\mathrm{FFI}}$” for the estimate (38), and “Aniso + NO” for the result at $\mathbf{k}^{\prime}=\mathbf{0}$ including the vacuum term (Sec. V.2).
Figure 13: Instability growth rate for some points identified on Fig. 12, under the different assumptions considered in this work (see text). Black circles represent a point that is unstable to the FFI, the red cross is unstable to slow modes, the blue triangle is unstable to the gapped CFI, and the orange star is unstable to the gapless CFI. The absence of a point means that the result was negative, hence showing no instability for that particular assumption.
The blue triangle and orange star points, located near the HMNS and close to the equatorial plane, show relatively little variation between the predicted growth rates. The blue triangle shows a gapped CFI, such that method A accurately predicts the isotropic CFI growth rate. The many-body corrections or scattering opacities do not make a significant change, just like the vacuum term (which is expected for a gapped mode, as discussed in Sec. IV.4). Including the anisotropies makes very little change, which is consistent given the small flux factors for this point (the largest flux factor is $\tilde{f}_{x}=0.035$, and the smallest is $\tilde{f}_{x}=0.004$). The dark orange point (which is also studied in Appendix C in comparison with a CCSN example) displays a gapless CFI, such that method A does not predict exactly the isotropic CFI growth rate, and as we showed in Sec. IV.1, method B typically overestimates the growth rate by a factor 2. The many-body corrections reduce the growth rate by a factor of a few, as expected from Fig. 7, bottom panel. Including the vacuum term slightly increases the growth rate, an effect that is enhanced in the anisotropic case (here the largest flux factor is $\tilde{f}_{x}=0.047$, and the smallest one is $\tilde{f}_{e}=0.011$).
@@
The final configuration we look at is located much further away from the HMNS, see the red cross on the left panel of Fig. 12. Focusing on the isotropic CFI, the reader may be surprised by the false predictions of methods A and B, when we showed in Sec. IV.1 that we could accurately estimate the multi-energy growth rate with those methods. However, looking at Fig. 3 one can see that the unstable energy-averaged modes shown on Fig. 13 are a gapless mode for method A, and a gapped one for method B. But our accurate estimate (35) combines gapped A modes and gapless B modes, which would give here, as expected, no instability. The inclusion of the vacuum term indicates a possible slow mode, but given the large distance of that point from the HMNS, neglecting the anisotropies is not justified ($\nu_{e}$ and $\bar{\nu}_{e}$ distributions have flux factors $\tilde{f}\simeq 0.7$). Indeed, we see that even though this point is not unstable to the FFI, the slow mode is enhanced by the anisotropies by about two orders of magnitude. This is representative of what happens in the region $X\geq 50\,\mathrm{km}$ on the left panel of Fig. 12. We leave for future work the detailed study of the mechanisms that lead to this anisotropy-driven growth rate enhancement, and their possible interplay with collisions.
@@ VI Conclusion
Linear stability analysis allows one to determine the locations and timescales of neutrino flavor instabilities, which correspond to an exponential growth of flavor coherence when one includes flavor mixing effects in neutrino transport. In this work, we have extended the angular moment-based LSA framework developed in [29] to describe a multi-energy system and included the vacuum and collision terms, which are respectively associated with slow and collisional instabilities. In order to avoid the inherent quantum closure problems associated with a moment method [115, 90, 91], we restricted the LSA to the “zero mode,” which is a wavevector coinciding with the homogeneous mode for isotropic background neutrino distributions, and for which a classical closure relation is sufficient. We have applied our LSA to a 7 ms postmerger snapshot from the “M1-NuLib” simulation presented in Ref. [52].
@@
Since we find many unstable modes at large distances from the HMNS, where neglecting the neutrino fluxes is not justified, we present in Sec. V a faithful assessment of flavor instabilities for the “zero,” $\mathbf{k}^{\prime}=\mathbf{0}$, mode, where the anisotropies of the neutrino field are taken into account. For the first time, we include not only absorption but also elastic scattering processes in the equations of motion in a multi-energy setting. To compensate for the underestimation of FFIs (which have the largest growth rates) due to the restricted wavenumber, we also include a growth rate estimate based on the depth of the FFI-associated angular crossing [Eq. (38)], which ensures that only slow and collisional modes can be missed. We find that the NSM snapshot we studied is mostly dominated by the FFI (see e.g., [51] for a discussion of the mechanisms leading to FFI in a similar system but with a long-lived HMNS), which in particular occurs in the regions where a resonancelike CFI could have taken place. Nevertheless, gapless and gapped CFIs still occur and extend the range of fast-flavor unstable regions. Finally, the slow modes at large distances appear to be enhanced by the anisotropies (compare Figs. 9 and 12, and see the red cross points in Fig. 13). Overall, Fig. 12 (and the bottom panels of Figs. 15 and 16 for other NSM snapshots) represents our most complete estimate of neutrino flavor instabilities in classical moment-based NSM simulation snapshots.
@@
As emphasized above, our moment LSA is restricted to the zero mode $\mathbf{k}^{\prime}=\mathbf{0}$, which avoids quantum closure issues and possible spurious modes at the expense of an incomplete assessment of the instabilities. Determining adequate closures for the generic $\mathbf{k}$ case, for instance based on multiangle LSA, is left for future work. It is also worth remembering that an assessment of flavor instabilities on a classically-computed simulation snapshot is not self-consistent, since instabilities would have modified the neutrino distributions in the simulation had they been included. An intermediate step towards an improved diagnosis of flavor instabilities in these environments could be to apply the framework presented here to simulations that already incorporate flavor conversion subgrid physics, such as [63, 116, 65, 64, 117]. Finally, we note that although highly useful, a LSA does not provide information on the final outcome of flavor instabilities. In particular for CFIs, although the post-instability asymptotic state has only been studied recently in simplified setups [66, 67], the smaller growth rates render almost unavoidable larger-scale simulations where advection and the time evolution of the flavor-diagonal components of the density matrices are taken into account. This is a challenging problem, for which moment approaches, with their inherently lower computational cost, present significant advantages.

### v2

@@ IV.3 Scattering processes and isotropy-breaking modes
where all the definitions can be found in Eq. (42); we do not reproduce them here for brevity. We assumed, without loss of generality, $\Delta_{\mathcal{N}}>\overline{\Delta}_{\mathcal{N}}$ (i.e., $A>0$), and we dropped the indices $\mathcal{N}$ and $\mathcal{F}$ as there is no difference here. Since $|\Delta_{\mathcal{N}}-\overline{\Delta}_{\mathcal{N}}|\geq\Delta_{\mathcal{N}}+\overline{\Delta}_{\mathcal{N}}$ (i.e., $|A|\leq G$), we see that the growth rate of the isotropy-breaking modes is smaller. However, the inclusion of isotropic scattering processes, which enter the flux equation but not the number density one [see Eq. (10)], can lead a priori to $\gamma_{\mathcal{F}}+(G\alpha_{\mathcal{F}})/A>-\gamma_{\mathcal{N}}+(G\alpha_{\mathcal{N}})/A$, if the inevitable increase in $\gamma_{\mathcal{F}}>\gamma_{\mathcal{N}}$ is compensated by an increase $\alpha_{\mathcal{F}}>\alpha_{\mathcal{N}}$.
@@ IV.4 Vacuum term and slow modes
Therefore, the imaginary part of Eq. (18) is not modified. However, in the regions with a gapless or no CFI, slow modes could play a role. We show the growth rates obtained from homogeneous and isotropic LSA in Fig. 9, top panel. They result from a combination of collisional and slow modes. In the bottom panel, we display the ratio of this growth rate with the default isotropic CFI-only (top panel of Fig. 2). As in Fig. 7, we show in turquoise the newly unstable regions, and in pink the regions which were unstable without the vacuum term but are stable with it. As expected from the argument above, the regions with the largest differences in growth rates correspond to the regions of gapless instability or of stability in Fig. 2. The results of Fig. 9 assume a normal ordering of neutrino masses—the results for the inverted ordering are very similar and we do not show them for brevity. The newly unstable regions are mostly associated with very small growth rates ($10^{4}\,\mathrm{s}^{-1}$), such that we do not expect a large impact of these slow modes. In particular, local studies of the slow instability in those regions would not be justified, since large-scale advection occurs on comparable or smaller timescales. Nevertheless, as we will show in the following, including the anisotropies of the neutrino field can lead to a significant enhancement of these growth rates (see Sec. V.2).
@@ V Results for anisotropic neutrino distributions
In this section and for the remainder of this paper, we no longer assume that neutrino distributions are isotropic, and we include the actual nonzero fluxes $\mathbf{F}_{\alpha\alpha}\neq\mathbf{0}$. This has two consequences. First, angular crossings between the neutrino and antineutrino distributions can occur, which are connected to fast instabilities [114, 115, 33]. Second, the separation between isotropy-preserving and isotropy-breaking modes—which can be done when the stability matrix has a block structure; see Eq. (40)—is not possible anymore. Furthermore, we also now consider different collision rates for number and flux, $\Gamma_{N}$ and $\Gamma_{F}$.
@@ V.1 Results at $k^{\prime}=0$, no vacuum term
Although our goal here is not to discuss in depth the origin of the angular crossings that lead to FFI, we note that our regions of FFI are consistent with other works. First, the instability patterns (small regions above and below the HMNS, extended regions in the tidal arms) are similar to the ones found in a 5 ms postmerger snapshot of another binary NSM simulation in [29] (see also Appendix D). Another comparison can be made with the recent detailed study [51], which used multiangle Boltzmann neutrino transport on top of fixed fluid backgrounds obtained from a $1.35M_{\odot}-1.35M_{\odot}$ NSM simulation leading to a long-lived ($1.3\,\mathrm{s}$) HMNS. While in the simulation we study the HMNS collapses to a black hole at $8.5\,\mathrm{ms}$ post-merger, some of the same physical mechanisms occur. For instance, the thin layers near the surface of the HMNS correspond to near-isotropic configurations with similar number densities of $\nu_{e}$ and $\bar{\nu}_{e}$, such that a small difference in flux between these two species leads to a crossing (see Sec. III.B.1. in [51]). Reference [51] points out other mechanisms, mostly based on the larger emission of $\bar{\nu}_{e}$ than $\nu_{e}$ at the surface of the HMNS, and the presence of an optically thick accretion disk with $\mu_{\nu}^{\mathrm{eq}}>0$. However, some of the processes leading to crossings are dependent on having sufficient angular resolution, and may be smeared out by using a two-moment scheme assuming maximum entropy angular distributions.
@@
Since the CFI is subdominant in the regions far from the HMNS, we focus upon the central region; see the right panel of Fig. 10. Contrary to what was observed in CCSNe [47], we find that the CFI is still largely present, extending the regions of instability. The growth rates are significantly smaller, but they could still lead to a buildup of flavor coherence throughout the equatorial region. The light-crossing time of a $1\,\mathrm{km}$ region is $\tau\simeq 1/(3\times 10^{5})\,\mathrm{s}$, which shows that instabilities with growth rates below a few times $10^{5}\,\mathrm{s}^{-1}$ can only have a cumulative, nonlocal effect (contrary to the FFI), which remains to be studied. We also caution that our LSA assumes constant and homogeneous flavor on-diagonal densities, such that our results only give us information on the local properties of the instabilities. Although this is beyond the scope of this work, we refer to [45] for a discussion of a possible way to perform a nonlocal LSA which takes into account the large-scale inhomogeneities of the classical neutrino field.
V.2 Fast, collisional and slow modes: Toward the full picture
@@
Note we have assumed $f_{\nu_{x}}=f_{\bar{\nu}_{x}}$. In these expressions, we have used the ansatz (22) for the neutrino spectra, where the angular part corresponds to maximum entropy distributions (23). Although the formula (38) may not be quantitatively exact, it ensures a FFI growth rate of order $\sqrt{2}G_{F}\mathcal{N}\sim 10^{10}\,\mathrm{s}^{-1}$—much larger than the CFI growth rates—for all regions where a crossing should be present (assuming maximum entropy angular distributions). We thus superimpose this estimate on the results at $\mathbf{k}^{\prime}=\mathbf{0}$, see Fig. 11. In the left panel, we see that now almost all CFI regions at large distances from the HMNS are overwhelmed by a FFI, but the enlarged version (right panel) still shows an extension of the instability regions in thin layers near the surface of the HMNS and, more significantly, in the tidal arms. We note that a detailed comparison of the FFI regions between Figs. 10 and 11 shows that the expression (38) can underestimate the FFI growth rate since the $\mathbf{k}^{\prime}=\mathbf{0}$ mode does not necessarily have the largest growth rate.
Beyond the quantitative uncertainty of Eq. (38), the Minerbo closure is an imperfect proxy for neutrino angular distributions. It was notably shown in Ref. [26] that, in NSM environments, instability regions vary depending on the reconstruction method used compared to benchmark multiangle Monte-Carlo calculations. Likewise, the use of multiangle Boltzmann transport in the CCSN simulations of Akaho et al. [65] allowed them to highlight the errors incurred with a Minerbo reconstruction for the detection of angular crossings. Despite these limitations, our use of moment-based NSM simulation data necessitates a closure-driven reconstruction. We note that, in the same situation but for CCSN simulations, Wang and Burrows [64] compared the few closure relations available which provide a full angular distribution. They found very similar CCSN dynamics when different angular reconstruction methods were used in a FFI subgrid model.
@@ V.3 Summary: Focus on a few example locations
In order to summarize the several growth rate determinations we have obtained in this work, and how different regions of a NSM simulation are subject to different flavor instabilities, we focus here on four locations identified in Fig. 12. The instability growth rate for these four points is shown in Fig. 13, where the marker matches the one used in Fig. 12 and each column corresponds to a different LSA assumption. From left to right, first assuming isotropic neutrino background distributions, “Iso CFI” is the multi-energy CFI, compared with the energy-averaged “Method A” (19) and “Method B” (21) (see Sec. IV.1), and with various multi-energy extensions: with many-body corrections to the absorption opacities (“Iso CFI, m-b,” Sec. IV.2), including scattering opacities (“Iso CFI + scatt,” Sec. IV.3), or with the vacuum term (“Iso CFI + NO,” Sec. IV.4). Then, we show the results obtained when we take into account the nonzero neutrino fluxes: “Aniso ($\mathbf{k}^{\prime}=\mathbf{0}$)” when the vacuum term is not included (Sec. V.1), “Estimate $\sigma_{\mathrm{FFI}}$” for the estimate (38), and “Aniso + NO” for the result at $\mathbf{k}^{\prime}=\mathbf{0}$ including the vacuum term (Sec. V.2).
Figure 13: Instability growth rate for some points identified in Fig. 12, under the different assumptions considered in this work (see text). Black circles represent a point that is unstable to the FFI, the red cross is unstable to slow modes, the blue triangle is unstable to the gapped CFI, and the orange star is unstable to the gapless CFI. The absence of a point means that the result was negative, hence showing no instability for that particular assumption.
The blue triangle and orange star points, located near the HMNS and close to the equatorial plane, show relatively little variation between the predicted growth rates. The blue triangle shows a gapped CFI, such that method A accurately predicts the isotropic CFI growth rate. The many-body corrections or scattering opacities do not make a significant change, just like the vacuum term (which is expected for a gapped mode, as discussed in Sec. IV.4). Including the anisotropies makes very little change, which is consistent given the small flux factors for this point (the largest flux factor is $\tilde{f}_{x}=0.035$, and the smallest is $\tilde{f}_{x}=0.004$). The dark orange point (which is also studied in Appendix C in comparison with a CCSN example) displays a gapless CFI, such that method A does not predict exactly the isotropic CFI growth rate, and as we showed in Sec. IV.1, method B typically overestimates the growth rate by a factor 2. The many-body corrections reduce the growth rate by a factor of a few, as expected from Fig. 7, bottom panel. Including the vacuum term slightly increases the growth rate, an effect that is enhanced in the anisotropic case (here the flux factors range from $\tilde{f}_{e}=0.011$ to $\tilde{f}_{x}=0.047$).
@@
The final configuration we look at is located much further away from the HMNS; see the red cross in the left panel of Fig. 12. Focusing on the isotropic CFI, the reader may be surprised by the false predictions of methods A and B, when we showed in Sec. IV.1 that we could accurately estimate the multi-energy growth rate with those methods. However, looking at Fig. 3 one can see that the unstable energy-averaged modes shown in Fig. 13 are a gapless mode for method A and a gapped one for method B. But our accurate estimate (35) combines gapped A modes and gapless B modes, which would give here, as expected, no instability. The inclusion of the vacuum term indicates a possible slow mode, but given the large distance of that point from the HMNS, neglecting the anisotropies is not justified ($\nu_{e}$ and $\bar{\nu}_{e}$ distributions have flux factors $\tilde{f}\simeq 0.7$). Indeed, we see that even though this point is not unstable to the FFI, the slow mode is enhanced by the anisotropies by about 2 orders of magnitude. This is representative of what happens in the region $X\geq 50\,\mathrm{km}$ on the left panel of Fig. 12. We leave for future work the detailed study of the mechanisms that lead to this anisotropy-driven growth rate enhancement, and their possible interplay with collisions.
@@ VI Conclusion
Linear stability analysis allows one to determine the locations and timescales of neutrino flavor instabilities, which correspond to an exponential growth of flavor coherence when one includes flavor mixing effects in neutrino transport. In this work, we have extended the angular moment-based LSA framework developed in [29] to describe a multi-energy system and included the vacuum and collision terms, which are, respectively, associated with slow and collisional instabilities. In order to avoid the inherent quantum closure problems associated with a moment method [116, 91, 92, 117], we restricted the LSA to the “zero mode,” which is a wavevector coinciding with the homogeneous mode for isotropic background neutrino distributions and for which a classical closure relation is sufficient. We have applied our LSA to a 7 ms postmerger snapshot from the “M1-NuLib” simulation presented in Ref. [52].
@@
Since we find many unstable modes at large distances from the HMNS, where neglecting the neutrino fluxes is not justified, we present in Sec. V a faithful assessment of flavor instabilities for the “zero,” $\mathbf{k}^{\prime}=\mathbf{0}$, mode, where the anisotropies of the neutrino field are taken into account. For the first time, we include not only absorption but also elastic scattering processes in the equations of motion in a multi-energy setting. To compensate for the underestimation of FFIs (which have the largest growth rates) due to the restricted wavenumber, we also include a growth rate estimate based on the depth of the FFI-associated angular crossing [Eq. (38)], which ensures that only slow and collisional modes can be missed. We find that the NSM snapshot we studied is mostly dominated by the FFI (see e.g., [51] for a discussion of the mechanisms leading to FFI in a similar system but with a long-lived HMNS), which in particular occurs in the regions where a resonancelike CFI could have taken place. Nevertheless, gapless and gapped CFIs still occur and extend the range of fast-flavor unstable regions. Finally, the slow modes at large distances appear to be enhanced by the anisotropies (compare Figs. 9 and 12, and see the red cross points in Fig. 13). Overall, Fig. 12 (and the bottom panels of Figs. 15 and 16 for other NSM snapshots) represents our most complete estimate of neutrino flavor instabilities in classical moment-based NSM simulation snapshots. In addition to flavor instabilities, other flavor conversion mechanisms can occur in NSM environments, such as MSW, turbulence-driven or matter-neutrino resonances [14]. For instance, the latter arise from the cancellation of the matter and self-interaction mean-field potentials [118, 119, 120, 121, 122, 123, 124, 125] and could have significant effects on nucleosynthesis yields [118, 126]. However, probing this mechanism requires tracking the propagation of neutrinos through the spatially varying potentials, and the interplay between these phenomena and flavor instabilities remains to be investigated.
@@
As emphasized above, our moment LSA is restricted to the zero mode $\mathbf{k}^{\prime}=\mathbf{0}$, which avoids quantum closure issues and possible spurious modes at the expense of an incomplete assessment of the instabilities. Determining adequate closures for the generic $\mathbf{k}$ case, for instance based on multiangle LSA, is left for future work. It is also worth remembering that an assessment of flavor instabilities on a classically computed simulation snapshot is not self-consistent, since instabilities would have modified the neutrino distributions in the simulation had they been included. An intermediate step toward an improved diagnosis of flavor instabilities in these environments could be to apply the framework presented here to simulations that already incorporate flavor conversion subgrid physics, such as [63, 127, 66, 64, 128]. Finally, we note that although highly useful, a LSA does not provide information on the final outcome of flavor instabilities. In particular for CFIs, although the post-instability asymptotic state has only been studied recently in simplified setups [67, 68], the smaller growth rates render almost unavoidable larger-scale simulations where advection and the time evolution of the flavor-diagonal components of the density matrices are taken into account. This is a challenging problem, for which moment approaches, with their inherently lower computational cost, present significant advantages.